# A Fourier-Jacobi Dirichlet series attached to modular forms of $\textup{SO}(2,4)$

### v1

@@ 5. Equivalence of Quadratic Spaces
with $A,B,C,D\in M_{2}(\mathcal{O}_{K})$. From [Wer19, Theorem 3.13], we must have $\det A=\det D=1$ and $f_{\omega}(A^{\sharp}B)=f_{\omega}(AC^{\sharp})=f_{\omega}(BD^{\sharp})=f_{\omega}(C^{\sharp}D)=0$, where denotes the adjugate of a matrix (see [Wer19, (3.9)]). But $f_{\omega}$ is injective and so $B=C=0$ and then $\widetilde{\gamma}=\epsilon(A)$, with $A\in\textup{SL}_{2}(\mathcal{O}_{K})$. Therefore, $\Psi$ is also surjective and hence an isomorphism.
Let now $u\in U^{+}$. We then have that $iu\in\mathbb{H}_{2}$. From [Wer19, Theorem 3.5], we have $f_{\omega}(M\langle Z\rangle)=\Phi(M)\langle f_{\omega}(Z)\rangle$ for all $Z\in\mathbb{H}_{2}$ and $M\in\textup{SU}(2,2;\mathcal{O}_{K})$. By taking now $M=\epsilon(\gamma)$ with $\gamma\in\textup{SL}_{2}(\mathcal{O}_{K})$ and using that $\Phi(\epsilon(\gamma))=\eta(\Psi(\gamma))$, (23) follows by unwinding the definitions of the actions.∎
We end this Section with the following identification of the orthogonal and Clifford groups of $(U,\psi)$, as is given by Shimura in [Shi06b, Proposition 2.2].
@@
Let $\mathcal{G}:=\{\alpha\in\textup{GL}_{2}(K)\mid\det(\alpha)\in\mathbb{Q}^{\times}\}$. There is an isomorphism $\Xi$ of $G^{+}(U)$ onto $\mathcal{G}$, such that if $\alpha\in G^{+}(U)$ and $\xi=\Xi(\alpha)$, then $\nu(\alpha)=\det(\xi)$ and $\tau(\alpha)h=\det(\xi)^{-1}\xi h\overline{\xi}^{t}$ for all $h\in U$. In particular, $\textup{SO}^{\psi}(U)$ is isomorphic to $\mathcal{G}/\mathbb{Q}^{\times}$.
6. The Dirichlet Series $D_{F,\xi_{i},\mathcal{P}}(s)$
In this Section, we will relate $D_{F,\xi_{i},\mathcal{P}}(s)$ of Proposition 4.2 to $L$-functions. We will first establish Lemma 4.1, namely that $\widehat{\Gamma}(L_{0})$ acts on $L_{0}\left[\ell,2^{-1}\mathbb{Z}\right]$ and there is a finite number of orbits under this action. For $\ell\in\mathbb{Z}_{>0}$, we define
@@ 6. The Dirichlet Series $D_{F,\xi_{i},\mathcal{P}}(s)$
The quotient $\Lambda^{+}\left[\ell,2^{-1}\mathbb{Z}\right]/\textup{SL}_{2}(\mathcal{O}_{K})$ is studied in detail by Shimura in [Shi06b]. In particular, we consider $\xi:=\textup{diag}(1,\ell)\in\Lambda^{+}\left[\ell,2^{-1}\mathbb{Z}\right]$. We define
$W:=\{u\in U\,\,|\,\,\psi(u,\xi)=0\},$ (24)
@@
We will now introduce localised and adelised groups and lattices, as usual. The reader can refer to [Shi97, Chapter 8] to recall the necessary definitions.
@@
$C_{p}:=\{g\in\textup{SO}^{\psi}(U)_{p}\,\,|\,\,g\Lambda_{p}=\Lambda_{p}\},$ (25)
$J_{p}:=\{\gamma\in G^{+}(U)_{p}\,\,|\,\,\tau(\gamma)\in C_{p},\,\,\nu(\gamma)\in\mathbb{Z}_{p}^{\times}\},$ (26)
@@
Shimura’s work in [Shi06b] gives the following Proposition.
@@
First of all, there is an equivalence of the sets $\Lambda^{+}\left[\ell,2^{-1}\mathbb{Z}\right]/\textup{SL}_{2}(\mathcal{O}_{K})$ and $\Lambda\left[\ell,2^{-1}\mathbb{Z}\right]/\Delta$, where $\Delta:=\{\alpha\in\textup{GL}_{2}(\mathcal{O}_{K})\mid\det(\alpha)=\pm 1\}$, as is shown in [Shi06b, p. 500]. Moreover, from the proof of [Shi06b, Theorem 2.10], the set $\Lambda[\ell,2^{-1}\mathbb{Z}]/\Delta$ is in bijection with $G^{+}(W)_{\mathbb{Q}}\backslash G^{+}(W)_{\mathbb{A}}/(G^{+}(W)_{\mathbb{A}}\cap J)$, because the class number of $\mathbb{Q}$ is $1$. ∎
@@
In order to now understand the Dirichlet series $D_{F,\xi_{i},\mathcal{P}}(s)$, we will work back in the quadratic space $(V_{0},\phi_{0})$. Note that $f_{\omega}$ of Lemma 5.1 induces an isomorphism between the Clifford algebras $A(U)$ and $A(V_{0})$, which then gives an isomorphism of the Clifford groups $G^{+}(U)$ and $G^{+}(V_{0})$. Moreover, $g_{\omega}$ of (19) is naturally extended to an isomorphism between $\textup{SO}^{\psi}(U)_{\mathbb{A}}$ and $\textup{SO}^{\phi_{0}}(V_{0})_{\mathbb{A}}$. We set
@@
Here, $\iota$ maps $\widehat{\Gamma}(L_{0})v\longmapsto\Gamma(L_{0})v$ for all $v\in L_{0}[\ell,2^{-1}\mathbb{Z}]$ and $\kappa,\lambda$ are the bijections defined by Shimura in [Shi06b, Theorem 1.6, (ii)] and [Shi06a, Theorem 2.2, (ii)], respectively.
@@
From [Shi06b, Theorem 1.6, (ii)], given $\epsilon\in G^{+}(\widetilde{W})_{\mathbb{A}}$, we take $\alpha\in G^{+}(V_{0})$ such that $\epsilon\in\alpha\widetilde{J}$ and set $\kappa(\epsilon):=\widehat{\Gamma}(L_{0})\tau(\alpha)^{-1}\xi$.
Similarly, from [Shi06a, Theorem 2.2, (ii)], given $\epsilon^{\prime}\in\widetilde{H}(\xi)_{\mathbb{A}}$, we take $\alpha^{\prime}\in\textup{SO}^{\phi_{0}}(V_{0})$ such that $\epsilon^{\prime}\in\alpha^{\prime}\widetilde{C}$ and set $\lambda(\epsilon^{\prime}):=\Gamma(L_{0})(\alpha^{\prime})^{-1}\xi$.
@@
We are now ready to deal with $D_{F,\xi_{i},\mathcal{P}}(s)$. This is done in [Psy25, Section 8] but we need to modify it slightly.
We observe that for each $i=1,\cdots,h_{1}$, the series $D_{F,\xi_{i},\mathcal{P}}(s)$ depends only on the $\Gamma^{+}(L_{0})$-equivalence class of $\xi_{i}$ (because of the invariance of $F$ under $\Gamma_{S}$, see also [Sch22, p. 26]). Since $\widehat{\Gamma}^{+}(L_{0})\leq\Gamma^{+}(L_{0})$, we may think the map $\iota$ of Proposition 6.2 as
@@
Therefore, we can write (15) as
$\mathcal{D}_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s)=\sum_{i=1}^{h_{1}}\zeta_{\xi_{i},\mathcal{P}}(s-k+3)D_{F,\iota(\xi_{i}),\mathcal{P}}(s).$ (27)
@@
$D_{F,\iota(\xi_{i}),\mathcal{P}}(s)=\zeta_{\mathcal{P}}\left(2s-2k+4\right)^{-1}L_{\mathcal{P}}\left(F;s-k+2\right)\times\\ \times\sum_{j=1}^{h_{2}}f_{j}(\lambda^{-1}(\iota(\xi_{i})))L_{\mathcal{P}}\left(\overline{f_{j}};s-k+5/2\right)^{-1}A_{f_{j}},$ (29)
using the notation of Proposition 6.2. This is [Psy25, (16)] tailored to our situation.
@@
$D_{F,\iota(\xi_{i}),\mathcal{P}}(s)=\zeta_{\mathcal{P}}\left(2s-2k+4\right)^{-1}L_{\mathcal{P}}\left(F;s-k+2\right)\times\\ \times\sum_{j=1}^{h_{2}}f_{j}(\tau(\zeta_{i}))L_{\mathcal{P}}\left(\overline{f_{j}};s-k+5/2\right)^{-1}A_{f_{j}}.$ (30)
@@ 7. Quaternion Algebras
Now that we have an understanding of the Dirichlet series $D_{F,\iota(\xi_{i}),\mathcal{P}}(s)$, our next step is to understand the zeta functions $\zeta_{\xi_{i}}(s)$ of (14). Our goal is to relate them to a partial zeta function in a quaternion algebra. To that end, we need several preparations. The quadratic space $(W,\psi)$ is three-dimensional and, as a result, its even Clifford algebra can be identified with a quaternion algebra, which in our case is a division algebra. Before making this explicit, let us recall some basic facts on quaternion algebras. We mainly follow [Voi21].
@@
These are orders in $B$ ([Voi21, Lemma 10.2.7]), and are called the left and right order of $I$, respectively. We say $I$ is right-invertible if there exists lattice $I^{\prime}\subset B$ such that $II^{\prime}=\mathcal{O}_{L}(I)$. Similarly, we define the notion of a left-invertible lattice, and we say $I$ is invertible if there is a two-sided inverse $I^{\prime}\subset B$, so that
@@
Now, if $\alpha\in B$, we define its reduced norm as $\textup{nrd}(\alpha):=\alpha\overline{\alpha}$. We define the reduced norm $\textup{nrd}(I)$ of a $\mathbb{Z}$-lattice $I$, as the $\mathbb{Z}$-submodule of $\mathbb{Q}$ generated by $\{\textup{nrd}(\alpha)\mid\alpha\in I\}$. This is fractional ideal of $\mathbb{Q}$ ([Voi21, Lemma 16.3.2]).
@@
Two $\mathbb{Z}$-lattices $I$ and $J$ are said to be in the same right class and write $I\sim_{R}J$, if $\exists\alpha\in B^{\times}$ such that $\alpha I=J$; similarly for the left. The relation $\sim_{R}$ defines an equivalence relation on the set of $\mathbb{Z}$-lattices in $B$ and we denote the equivalence class of a $\mathbb{Z}$-lattice $I$ by $[I]_{R}$. The (right) class set of $\mathcal{O}$ is
@@
We say two orders $\mathcal{O},\mathcal{O^{\prime}}$ are of the same type, if $\exists\alpha\in B^{\times}$ such that $\mathcal{O}^{\prime}=\alpha^{-1}\mathcal{O}\alpha$. Note that this is true if and only if $\mathcal{O}$ and $\mathcal{O}^{\prime}$ are isomorphic as $\mathbb{Z}$-algebras. Moreover, we say that $\mathcal{O}$ and $\mathcal{O}^{\prime}$ are locally isomorphic, if the localisations $\mathcal{O}_{p}$ and $\mathcal{O}^{\prime}_{p}$ are of the same type for all rational primes $p$. Here, we define the localisation of a $\mathbb{Z}$-lattice in $B$ as usual (see, for example, [Voi21, Section 9.4]).
@@
Let us now define the quaternion algebra $B:=K+\epsilon K$ with $\epsilon^{2}=$-$\ell$ and $\epsilon a=\overline{a}\epsilon$ for all $a\in K$. Let also $\mathfrak{o}:=\mathcal{O}_{K}+\epsilon\mathcal{O}_{K}$ denote an order in $B$. From the proof of [Shi06b, Theorem 2.10], we have $\Xi(A^{+}(W))=B$ and $\Xi(G^{+}(W))=B^{\times}$, where $\Xi$ is the map of Proposition 5.4. Moreover, from the same Theorem, $\Xi(G^{+}(W)_{p}\cap J_{p})=\mathfrak{o}_{p}^{\times}$ for all $p$.
@@
Because the lattice $L_{0}$ is taken maximal in $(V_{0},\phi_{0})$, we have from [Shi06, Paragraph 6.3] that $[L_{0}^{*}/L_{0}]=\det(S_{0})\mathbb{Z}$. Moreover, as in [Psy25, Lemma 9.1], $\widetilde{H}(\xi)$ is represented via the matrix
@@
with respect to a basis of $M:=\mathbb{Z}^{3}$. We will now show that under our assumptions on $k$ and $m$, $M$ is a maximal $\mathbb{Z}$–lattice.
If $m\equiv 3\pmod{4}$, then $\det(2T)=2\ell\det(S)=2\ell m$ and because of the conditions on $\ell$ and $m$, $\det(2T)/2$ is square-free. Hence, from [Sch22, Proposition 1.6.12], the lattice $\mathbb{Z}^{3}$ is maximal with respect to $(\widetilde{W},\phi_{0})$.
If $m\equiv 1\pmod{4}$, we will use [Sch22, Lemma 1.6.5 (ii)]. Assume we pick $x=(x_{1},x_{2},x_{3})\in M^{*}$ such that $T[x]\in\mathbb{Z}$. We will show that $x\in M$, which will give us that $M$ is $\mathbb{Z}$–maximal. We write $x_{1}=y_{1}/(2\ell),\ x_{2}=y_{2}/2,\ x_{3}=y_{3}/(2m)$, where $y_{i}\in\mathbb{Z}$. Then, the condition $T[x]\in\mathbb{Z}$ is equivalent to $4\ell m\mid my_{1}^{2}+\ell my_{2}^{2}+\ell y_{3}^{2}$. Now, by the assumptions, $\ell$ and $4m$ are coprime. Hence, we can look at each divisibility separately.
@@
Now, again from [Shi06, Paragraph 6.3], we have $[M^{*}/M]=\det(2T)\mathbb{Z}=2\ell\det(S)\mathbb{Z}$. Since $\phi_{0}[f_{\omega}(\xi)]=\ell$ and $2\phi_{0}(f_{\omega}(\xi),L_{0})=\mathbb{Z}$, we have $\phi_{0}[f_{\omega}(\xi)]\phi_{0}(f_{\omega}(\xi),L_{0})^{-2}=4\ell\mathbb{Z}$ and this equals $2[M^{*}/M][L_{0}^{*}/L_{0}]^{-1}$. Hence, from [Mur18, Corollary 4.3], we obtain $L_{0}\cap\widetilde{W}$ is $\mathbb{Z}$-maximal in $\widetilde{W}$.
This now means that $(L_{0}\cap\widetilde{W})_{p}$ is maximal in $\widetilde{W}_{p}$ for all primes $p$. Hence, from [Shi04, Proposition 11.12, (iv)], we obtain $\widetilde{H}(\xi)_{p}\cap\widetilde{C}=D_{p}$ for all primes $p$, as $t_{p}$ is always even in our case. Here, $D_{p}:=D\cap\widetilde{H}(\xi)_{p}$. Hence, similarly to [Psy25, Proposition 9.3], we obtain $D=\widetilde{H}(\xi)_{\mathbb{A}}\cap\widetilde{C}$, as required. ∎
@@
We now have a surjective map (see [Voi21, Lemma 17.4.13])

### v2

@@ 7. Equivalence of Quadratic Spaces
with $A,B,C,D\in M_{2}(\mathcal{O}_{K})$. From [wernz, Theorem 3.13], we must have $\det A=\det D=1$ and $f_{\omega}(A^{\sharp}B)=f_{\omega}(AC^{\sharp})=f_{\omega}(BD^{\sharp})=f_{\omega}(C^{\sharp}D)=0$, where denotes the adjugate of a matrix (see [wernz, (3.9)]). But $f_{\omega}$ is injective and so $B=C=0$ and then $\widetilde{\gamma}=\epsilon(A)$, with $A\in\textup{SL}_{2}(\mathcal{O}_{K})$. Therefore, $\Psi$ is also surjective and hence an isomorphism.
Let now $u\in U^{+}$. We then have that $iu\in\mathbb{H}_{2}$. From [wernz, Theorem 3.5], we have $f_{\omega}(M\langle Z\rangle)=\Phi(M)\langle f_{\omega}(Z)\rangle$ for all $Z\in\mathbb{H}_{2}$ and $M\in\textup{SU}(2,2;\mathcal{O}_{K})$. By taking now $M=\epsilon(\gamma)$ with $\gamma\in\textup{SL}_{2}(\mathcal{O}_{K})$ and using that $\Phi(\epsilon(\gamma))=\eta(\Psi(\gamma))$, (7.6) follows by unwinding the definitions of the actions.∎
We end this Section with the following identification of the orthogonal and Clifford groups of $(U,\psi)$, as is given by Shimura in [shimura2006quadratic, Proposition 2.2].
@@
Let $\mathcal{G}:=\{\xi\in\textup{GL}_{2}(K)\mid\det(\xi)\in\mathbb{Q}^{\times}\}$. There is an isomorphism $\Xi$ of $G^{+}(U)$ onto $\mathcal{G}$, such that if $\alpha\in G^{+}(U)$ and $\xi=\Xi(\alpha)$, then $\nu(\alpha)=\det(\xi)$ and $\tau(\alpha)h=\det(\xi)^{-1}\xi h\overline{\xi}^{t}$ for all $h\in U$. In particular, $\textup{SO}^{\psi}(U)$ is isomorphic to $\mathcal{G}/\mathbb{Q}^{\times}$.
8. The Dirichlet Series $D_{\mathcal{Q}}(F,\xi_{i};s)$
In this Section, we will relate $D_{\mathcal{Q}}(F,\xi_{i};s)$ of (6.5) to $L$-functions. We will first establish Lemma 6.1, namely that $\widehat{\Gamma}(L_{0})$ acts on $L_{0}\left[\ell,2^{-1}\mathbb{Z}\right]$ and there is a finite number of orbits under this action. For $\ell\in\mathbb{Z}_{>0}$, we define
@@ 8. The Dirichlet Series $D_{\mathcal{Q}}(F,\xi_{i};s)$
The quotient $\Lambda^{+}\left[\ell,2^{-1}\mathbb{Z}\right]/\textup{SL}_{2}(\mathcal{O}_{K})$ is studied in detail by Shimura in [shimura2006quadratic]. In particular, we consider $\xi:=\textup{diag}(1,\ell)\in\Lambda^{+}\left[\ell,2^{-1}\mathbb{Z}\right]$. We define
(8.1) $W:=\{u\in U\,\,|\,\,\psi(u,\xi)=0\},$
@@
We will now introduce localised and adelised groups and lattices, as usual. The reader can refer to [Shimura_Euler_Product, Chapter 8] to recall the necessary definitions.
@@
(8.2) $C_{p}:=\{g\in\textup{SO}^{\psi}(U)_{p}\,\,|\,\,g\Lambda_{p}=\Lambda_{p}\},$
(8.3) $J_{p}:=\{\gamma\in G^{+}(U)_{p}\,\,|\,\,\tau(\gamma)\in C_{p},\,\,\nu(\gamma)\in\mathbb{Z}_{p}^{\times}\},$
@@
Shimura’s work in [shimura2006quadratic] gives the following Proposition.
@@
First of all, there is an equivalence of the sets $\Lambda^{+}\left[\ell,2^{-1}\mathbb{Z}\right]/\textup{SL}_{2}(\mathcal{O}_{K})$ and $\Lambda\left[\ell,2^{-1}\mathbb{Z}\right]/\Delta$, where $\Delta:=\{\alpha\in\textup{GL}_{2}(\mathcal{O}_{K})\mid\det(\alpha)=\pm 1\}$, as is shown in [shimura2006quadratic, p. 500]. Moreover, from the proof of [shimura2006quadratic, Theorem 2.10], the set $\Lambda[\ell,2^{-1}\mathbb{Z}]/\Delta$ is in bijection with $G^{+}(W)_{\mathbb{Q}}\backslash G^{+}(W)_{\mathbb{A}}/(G^{+}(W)_{\mathbb{A}}\cap J)$, because the class number of $\mathbb{Q}$ is $1$. ∎
@@
In order to now understand the Dirichlet series $D_{F,\xi_{i},\mathcal{P}}(s)$, we will work back in the quadratic space $(V_{0},\phi_{0})$. Note that $f_{\omega}$ of Lemma 7.1 induces an isomorphism between the Clifford algebras $A(U)$ and $A(V_{0})$, which then gives an isomorphism of the Clifford groups $G^{+}(U)$ and $G^{+}(V_{0})$. Moreover, $g_{\omega}$ of (7.2) is naturally extended to an isomorphism between $\textup{SO}^{\psi}(U)_{\mathbb{A}}$ and $\textup{SO}^{\phi_{0}}(V_{0})_{\mathbb{A}}$. We set
@@
Here, $\iota$ maps $\widehat{\Gamma}(L_{0})v\longmapsto\Gamma(L_{0})v$ for all $v\in L_{0}[\ell,2^{-1}\mathbb{Z}]$ and $\kappa,\lambda$ are the bijections defined by Shimura in [shimura2006quadratic, Theorem 1.6, (ii)] and [shimura_diophantine_2006, Theorem 2.2, (ii)], respectively.
@@
From [shimura2006quadratic, Theorem 1.6, (ii)], given $\epsilon\in G^{+}(\widetilde{W})_{\mathbb{A}}$, we take $\alpha\in G^{+}(V_{0})$ such that $\epsilon\in\alpha\widetilde{J}$ and set $\kappa(\epsilon):=\widehat{\Gamma}(L_{0})\tau(\alpha)^{-1}\xi$.
Similarly, from [shimura_diophantine_2006, Theorem 2.2, (ii)], given $\epsilon^{\prime}\in\widetilde{H}(\xi)_{\mathbb{A}}$, we take $\alpha^{\prime}\in\textup{SO}^{\phi_{0}}(V_{0})$ such that $\epsilon^{\prime}\in\alpha^{\prime}\widetilde{C}$ and set $\lambda(\epsilon^{\prime}):=\Gamma(L_{0})(\alpha^{\prime})^{-1}\xi$.
@@
We are now ready to deal with $D_{\mathcal{Q}}(F,\xi_{i};s)$. This is done in [psyroukis_orthogonal, Section 8] but we need to modify it slightly.
We observe that for each $i=1,\cdots,h_{1}$, the series $D_{\mathcal{Q}}(F,\xi_{i};s)$ depends only on the $\Gamma^{+}(L_{0})$-equivalence class of $\xi_{i}$ (because of the invariance of $F$ under $\Gamma_{S}$, see also [eisenstein_thesis, p. 26]). Since $\widehat{\Gamma}^{+}(L_{0})\leq\Gamma^{+}(L_{0})$, we may think the map $\iota$ of Proposition 8.2 as
@@
Therefore, we can write $D(\bm{F},\bm{P};s)$ as
(8.4) $D(\bm{F},\bm{P};s)=\sum_{i=1}^{h_{1}}\left(\sum_{\mathfrak{m}\in\textup{Cl}(K)}\zeta_{\mathfrak{m},\mathcal{Q}}(s-k+3)\zeta_{\mathcal{Q}}(\xi_{\mathfrak{m},i};s-k+3)\right)D_{\mathcal{Q}}(F,\iota(\xi_{i});s),$
@@
(8.6) $D_{\mathcal{Q}}(F,\iota(\xi_{i});s)=\zeta_{\mathcal{Q}}\left(2s-2k+4\right)^{-1}L_{\mathcal{Q}}\left(F;s-k+2\right)\times\\ \times\sum_{j=1}^{h_{2}}f_{j}(\lambda^{-1}(\iota(\xi_{i})))L_{\mathcal{Q}}\left(\overline{f_{j}};s-k+5/2\right)^{-1}A_{f_{j}},$
using the notation of Proposition 8.2. This is [psyroukis_orthogonal, (16)] tailored to our situation.
@@
(8.7) $D_{\mathcal{Q}}(F,\iota(\xi_{i});s)=\zeta_{\mathcal{Q}}\left(2s-2k+4\right)^{-1}L_{\mathcal{Q}}\left(F;s-k+2\right)\times\\ \times\sum_{j=1}^{h_{2}}f_{j}(\tau(\zeta_{i}))L_{\mathcal{Q}}\left(\overline{f_{j}};s-k+5/2\right)^{-1}A_{f_{j}}.$
@@ 9. Quaternion Algebras
Now that we have an understanding of the Dirichlet series $D_{\mathcal{Q}}(F,\iota(\xi_{i});s)$, our next step is to understand the zeta functions $\zeta_{\mathcal{Q}}(\xi_{\mathfrak{m},i};s)$ of (6.4). Our goal is to relate them to a partial zeta function in a quaternion algebra. To that end, we need several preparations. The quadratic space $(W,\psi)$ is three-dimensional and, as a result, its even Clifford algebra can be identified with a quaternion algebra, which in our case is a division algebra. Before making this explicit, let us recall some basic facts on quaternion algebras. We mainly follow [voight].
@@
These are orders in $B$ ([voight, Lemma 10.2.7]), and are called the left and right order of $I$, respectively. We say $I$ is right-invertible if there exists a lattice $I^{\prime}\subset B$ such that $II^{\prime}=\mathcal{O}_{L}(I)$. Similarly, we define the notion of a left-invertible lattice, and we say $I$ is invertible if there is a two-sided inverse $I^{\prime}\subset B$, so that
@@
Now, if $\alpha\in B$, we define its reduced norm as $\textup{nrd}(\alpha):=\alpha\overline{\alpha}$. We define the reduced norm $\textup{nrd}(I)$ of a $\mathbb{Z}$-lattice $I$, as the $\mathbb{Z}$-submodule of $\mathbb{Q}$ generated by $\{\textup{nrd}(\alpha)\mid\alpha\in I\}$. This is a fractional ideal of $\mathbb{Q}$ ([voight, Lemma 16.3.2]).
@@
Two $\mathbb{Z}$-lattices $I$ and $J$ are said to be in the same right class and write $I\sim_{R}J$, if $\exists\alpha\in B^{\times}$ such that $\alpha I=J$; similarly for the left. The relation $\sim_{R}$ defines an equivalence relation on the set of $\mathbb{Z}$-lattices in $B$, and we denote the equivalence class of a $\mathbb{Z}$-lattice $I$ by $[I]_{R}$. The (right) class set of $\mathcal{O}$ is
@@
We say two orders $\mathcal{O},\mathcal{O^{\prime}}$ are of the same type, if $\exists\alpha\in B^{\times}$ such that $\mathcal{O}^{\prime}=\alpha^{-1}\mathcal{O}\alpha$. Note that this is true if and only if $\mathcal{O}$ and $\mathcal{O}^{\prime}$ are isomorphic as $\mathbb{Z}$-algebras. Moreover, we say that $\mathcal{O}$ and $\mathcal{O}^{\prime}$ are locally isomorphic if the localisations $\mathcal{O}_{p}$ and $\mathcal{O}^{\prime}_{p}$ are of the same type for all rational primes $p$. Here, we define the localisation of a $\mathbb{Z}$-lattice in $B$ as usual (see, for example, [voight, Section 9.4]).
@@
Let us now define the quaternion algebra $B:=K+\epsilon K$ with $\epsilon^{2}=$-$\ell$ and $\epsilon a=\overline{a}\epsilon$ for all $a\in K$. Let also $\mathfrak{o}:=\mathcal{O}_{K}+\epsilon\mathcal{O}_{K}$ denote an order in $B$. From the proof of [shimura2006quadratic, Theorem 2.10], we have $\Xi(A^{+}(W))=B$ and $\Xi(G^{+}(W))=B^{\times}$, where $\Xi$ is the map of Proposition 7.4. Moreover, from the same Theorem, $\Xi(G^{+}(W)_{p}\cap J_{p})=\mathfrak{o}_{p}^{\times}$ for all $p$.
@@
Because of Proposition 8.2, we also want to identify the the quotient $H(\xi)_{\mathbb{Q}}\backslash H(\xi)_{\mathbb{A}}/(H(\xi)_{\mathbb{A}}\cap C)$ in the quaternion algebra $B$. To this end, we will turn back to the quadratic space $(V_{0},\phi_{0})$. We have the following Proposition.
@@
Because the lattice $L_{0}$ is taken maximal in $(V_{0},\phi_{0})$, we have from [shimura_classification_quadratic_forms, Paragraph 6.3] that $[L_{0}^{*}/L_{0}]=\det(S_{0})\mathbb{Z}$. Moreover, as in [psyroukis_orthogonal, Lemma 9.1], $\widetilde{H}(\xi)$ is represented via the matrix
@@
with respect to a basis of $M:=\mathbb{Z}^{3}$. We will now show that under our assumptions on $\ell$ and $m$, $M$ is a maximal $\mathbb{Z}$–lattice.
If $m\equiv 3\pmod{4}$, then $\det(2T)=2\ell\det(S)=2\ell m$ and because of the conditions on $\ell$ and $m$, $\det(2T)/2$ is square-free. Hence, from [eisenstein_thesis, Proposition 1.6.12], the lattice $\mathbb{Z}^{3}$ is maximal with respect to $(\widetilde{W},\phi_{0})$.
If $m\equiv 1\pmod{4}$, we will use [eisenstein_thesis, Lemma 1.6.5 (ii)]. Assume we pick $x=(x_{1},x_{2},x_{3})\in M^{*}$ such that $T[x]\in\mathbb{Z}$. We will show that $x\in M$, which will give us that $M$ is $\mathbb{Z}$–maximal. We write $x_{1}=y_{1}/(2\ell),\ x_{2}=y_{2}/2,\ x_{3}=y_{3}/(2m)$, where $y_{i}\in\mathbb{Z}$. Then, the condition $T[x]\in\mathbb{Z}$ is equivalent to $4\ell m\mid my_{1}^{2}+\ell my_{2}^{2}+\ell y_{3}^{2}$. Now, by the assumptions, $\ell$ and $4m$ are coprime. Hence, we can look at each divisibility separately.
@@
Now, again from [shimura_classification_quadratic_forms, Paragraph 6.3], we have $[M^{*}/M]=\det(2T)\mathbb{Z}=2\ell\det(S)\mathbb{Z}$. Since $\phi_{0}[f_{\omega}(\xi)]=\ell$ and $2\phi_{0}(f_{\omega}(\xi),L_{0})=\mathbb{Z}$, we have $\phi_{0}[f_{\omega}(\xi)]\phi_{0}(f_{\omega}(\xi),L_{0})^{-2}=4\ell\mathbb{Z}$ and this equals $2[M^{*}/M][L_{0}^{*}/L_{0}]^{-1}$. Hence, from [murata_preprint, Corollary 4.3], we obtain $L_{0}\cap\widetilde{W}$ is $\mathbb{Z}$-maximal in $\widetilde{W}$.
This now means that $(L_{0}\cap\widetilde{W})_{p}$ is maximal in $\widetilde{W}_{p}$ for all primes $p$. Hence, from [shimura2004arithmetic, Proposition 11.12, (iv)], we obtain $\widetilde{H}(\xi)_{p}\cap\widetilde{C}=D_{p}$ for all primes $p$, as $t_{p}$ is always even in our case. Here, $D_{p}:=D\cap\widetilde{H}(\xi)_{p}$. Hence, similarly to [psyroukis_orthogonal, Proposition 9.3], we obtain $D=\widetilde{H}(\xi)_{\mathbb{A}}\cap\widetilde{C}$, as required. ∎
@@
We now have a surjective map (see [voight, Lemma 17.4.13])